OpenStudy (anonymous):

give the next three numbers in the pattern 1,3,7,15 so: 1-3? 3-7? and 15-7?

OpenStudy (jamesj):

better to look at 3 - 1 = 2 7 - 3 = 4 15 - 7 = ... ?

OpenStudy (anonymous):

8!!

OpenStudy (jamesj):

Go back to the questions you've already asked and make sure you know the answers. This one is half a step more difficult and if you don't understand the ones before this one, you will find this one difficult.

OpenStudy (anonymous):

oh i was not right..

OpenStudy (anonymous):

Wait i it 6!

OpenStudy (anonymous):

Is it*